Suppose that a perfectly competitive firm manufactures gizmos with the following cost structure (including all opportunity costs):

Quantity of gizmos Total Cost ($)
0 75
1 150
2 250
3 425
4 675

a. calculate the marginal cost schedule for this firm in a table, and then graph the marginal cost curve.

b. if the price of gizmos on the market is $175 each, how many gizmos should the firm produce to maximize profits? what is the level of the firm's revenues at its chosen output level? how much does it make in profit?

c. suppose that more firms start producing gizmos, and the market price drops to $125. how many gizmos should this firm produce to maximize profits? (note: in the case of discrete quantities such as these, interpret the P=MC rule as "produce as long as price is at least as great as marginal cost.") What is this firm's new revenue level? How much does it make in profits?

d. when the price is $125, will more firms want to enter the market? will existing firms want to exit?
Answer #1

a)

Q MC

0 -

1 75

2 100

3 175

4 250

following diagram:

MC Cost

b)

When price is $ 175, firm would be producing 3 units.

TR= 175*3

= 525

TC = 425

Profit = 525 - 425

= $ 100

c)

two units would be produced if price falls to $ 125

Profit = 125*2 -250

= 0

Firm is earning zero economic profit or normal profit

d)

There would be neither entry nor exit of firms. Since firms are earning only normal profits. or zero economic profit.